# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

### v1

@@ 6.8 Transformers and LLMs-based ASR
Similarly, the LM aspect of LLM based on DTL can enhance ASR by leveraging TL. Pre-training LLM on vast text corpora equips it with extensive language representations, aiding in addressing diverse language challenges. Fine-tuning enables adaptation to specific language characteristics, improving transcription accuracy and contextual appropriateness. Both AM and LM fine-tuning can benefit from DA, incorporating target domain data to tailor models, reducing domain mismatch, and enhancing effectiveness and generalization. Utilizing LLM for objective ASR testing and MOS evaluation involves compiling diverse datasets, fine-tuning LLM, and integrating it into the ASR system. Evaluation metrics like CER and Pearson’s correlation gauge system performance, guiding further fine-tuning iterations for improved results. This iterative process ensures the ASR system’s continual enhancement and accurate MOS scale generation. Researchers are invited to explore these gaps and advance the integration of more advanced AI techniques such as DRL and FL into both AM and LM domains within transformer-based models. Additionally, there is a need for further investigation into Transformers and DTL-based ASR schemes specifically tailored for the LM domain. Closing these gaps will contribute to the development of more robust and effective language models across various applications.
@@ 7 Conclusion
This survey has provided a comprehensive review of DTL, FL, and RL-based ASR frameworks, offering insights into the latest developments and helping researchers and professionals understand current challenges. Additionally, the integration of transformers, powerful DL models, has been explored for their ability to capture complex dependencies in ASR sequences. By presenting a structured taxonomy and conducting critical analyses, this paper has shed light on the strengths and weaknesses of existing frameworks, as well as highlighted ongoing challenges. Moving forward, further research is needed to overcome these challenges and unlock the full potential of advanced DL techniques in ASR. Future work should focus on refining existing approaches, addressing privacy concerns in FL, improving RL algorithms for ASR optimization, and exploring innovative ways to leverage transformers for more efficient and accurate speech recognition. By continuing to innovate and collaborate across disciplines, we can push the boundaries of ASR technology and realize its transformative impact on various fields, including healthcare, communication, and accessibility.

### v2

@@ 5.6 Transformers and LLMs-based ASR
Similarly, the LM aspect of LLM based on DTL can enhance ASR by leveraging DTL. Pre-training LLM on vast text corpora equips it with extensive language representations, aiding in addressing diverse language challenges. Fine-tuning enables adaptation to specific language characteristics, improving transcription accuracy and contextual appropriateness. Both AM and LM fine-tuning can benefit from DA, incorporating target domain data to tailor models, reducing domain mismatch, and enhancing effectiveness and generalization. Utilizing LLM for objective ASR testing and MOS evaluation involves compiling diverse datasets, fine-tuning LLM, and integrating it into the ASR system. Evaluation metrics like CER and Pearson’s correlation gauge system performance, guiding further fine-tuning iterations for improved results. This iterative process ensures the ASR system’s continual enhancement and accurate MOS scale generation. Researchers are invited to explore these gaps and advance the integration of more advanced AI techniques such as DRL and FL into both AM and LM domains within Transformer-based models. Additionally, there is a need for further investigation into Transformers and DTL-based ASR schemes specifically tailored for the LM domain. Closing these gaps will contribute to the development of more robust and effective language models across various applications.
@@ 6 Conclusion
This survey has provided a comprehensive review of DTL, FL, and DRL-based DTL frameworks, offering insights into the latest developments and helping researchers and professionals understand current challenges. Additionally, the integration of Transformers, powerful DL models, has been explored for their ability to capture complex dependencies in DTL sequences. By presenting a structured taxonomy and conducting critical analyses, this paper has shed light on the strengths and weaknesses of existing frameworks, as well as highlighted ongoing challenges. Moving forward, further research is needed to overcome these challenges and unlock the full potential of advanced DL techniques in DTL. Future work should focus on refining existing approaches, addressing privacy concerns in FL, improving DRL algorithms for DTL optimization, and exploring innovative ways to leverage Transformers for more efficient and accurate speech recognition. By continuing to innovate and collaborate across disciplines, we can push the boundaries of ASR technology and realize its transformative impact on various fields, including healthcare, communication, and accessibility.